OpenStudy (anonymous):

factor x^2-25 ------- x + 5 ?????

OpenStudy (anonymous):

(x+5)*(x-5)

OpenStudy (anonymous):

but once again, how did you get to that solution?

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

= (x-5)(x+5) / x+ 5 = x-5

OpenStudy (anonymous):

Notice how 25 is the square of 5, and the fact that there are no single power versions of x. That's a big hint when trying to foil exponentials. From there i simply work backwords and multiply x+5 and x-5 to check my guess. Works about 95% of the time.

OpenStudy (anonymous):

wow I don't think I will ever get the hang of this....thanks for your help!